Before the Christmas shopping rush began, a department store had noted that the percentage of customers who use a major credit card, the percentage who use the store’s credit card, the percentage who pay by debit card, and the percentage who pay with cash are all the same, 25% for each. During the Christmas season, in a sample of 200 shoppers, 68 used a major credit card, 42 used the store’s credit card, 45 paid with cash, and 45 paid with a debit card. At a 95% confidence level, test to see if the methods of payment have changed during the Christmas shopping rush. please show your work
Answer #1

Hypothesis :

H_{0}:P_{1}=P_{2}=P_{3}=P_{4}=0.25

H_{1}:P_{i}\neq 0.25 (at least one Pi is not equal to 0.25 )

Test statistics:

\chi ^{2}=\sum \frac{(Oi-Ei)^{2}}{Ei}

Oi =Observed frequency

Ei=Expcted frequency

K number of observation

Ei=200*0.25=50

Oi Ei (Oi-Ei)2 (Oi-Ei)2/Ei
1 68 50 324 6.48
2 42 50 64 1.28
3 45 50 25 0.5
4 45 50 25 0.5
Total 200 200 8.76

\chi ^{2}=8.76

Df=n-1=4-1=3

P value : Is the value of the lowets level of singnificance at which we could reject the Null hypothesis H0
Using P value approch ,we reject the null hypothesis

p value at 0.05 leve of significane

P value =0.0326<0.05

Conclusion : result is significant at 0.05 level of significant